OpenStudy (anonymous):

x/3+2x/15<2/3 Find an inequality

hero (hero):

I think you mean "Solve" instead of "Find"

OpenStudy (anonymous):

Yes

hero (hero):

Multiply the first fraction x/3 by 5/5. Let me know what you get.

OpenStudy (anonymous):

I am having a problem how to solve this

hero (hero):

Do you know how to multiply: \[\frac{x}{3} \times \frac{5}{5}\]